a ball player catches a ball 3.53 seconds after throwing it vertically upward.
- with what speed did he throw it?
-what height did it reach?
1
Expert's answer
2019-06-05T09:57:22-0400

Z coordinate of the ball depends on time as


"z(t) = v_0 t -\frac{gt^2}{2}"


where g is "9.8 \,\text{ms}^{-2}" - free fall acceleration., "v_0" is initial speed of the ball.

a ball player catches a ball "\tau = 3.53" seconds after throwing it, therefore


"z(\tau) = v_0 \tau -\frac{g\tau^2}{2}=0 \Rightarrow v_0= \frac{g\tau}{2} = 17.3\,\text{ms}^{-1}"

The Ball reached maximum heigth at mid time between throwing ball and catching it:


"H = z(\tau/2) = \frac{g \tau^2}{8}=15.3\,\text{m}"